# On the Baltimore Light RailLink into the quantum future

### v1

@@ III.4 Real-world railway rescheduling perspective
For this particular problem, the input disturbed timetable includes conflicts between three pairs of trains. A train diagram representing this timetable is shown in Fig. 10a. To resolve this conflict, it is necessary that some of the trains involved in conflicts add an additional wait to their schedule. Figure 10b shows an optimal rescheduling that resolves these conflicts, generated with a classical ILP solver. Notice that passing times of all the trains are the same (i.e., the slopes of the lines are identical). This is not what would be expected for a timetable describing the movement of trains through stochastic areas of the rail network; in such a case, the passing times should be randomly perturbed and so the slopes should vary. This feature is visible in Fig. 10c, which depict timetables corresponding to the best (lowest energy) solution returned by the D-Wave annealer and a solutions corresponding to a low-lying excited state, respectively. This observation supports the assertion that when NISQ devices are used to solve scheduling problems with this type of QUBO representation, their inherent noise may manifest in a way that can capture stochastic behavior of the underlying real-world model.
To elaborate this observation, we compare the statistics of the D-Wave output to the statistics measured from real-world rail traffic. For this comparison, we present in Fig. 11 the distribution of passing times northbound ${\rm MR}\rightarrow{\rm CS}$ and southbound ${\rm CS}\rightarrow{\rm MR}$. The data  [^5] Baltimore Light RailLink passing time data collected with Swiftly’s GTFS-realtime API (https://swiftly.zendesk.com/hc/en-us) and covers workdays from 11.1.24 - 31.1.24, excluding the 17th, 18th, and 26th due to issues with the Baltimore Light RailLink system. was collected by aggregating passing times at peak hour during work days, as that is when the random disturbances due to road traffic are most intense.
@@ IV Conclusions
At present, quantum computing is in the era of NISQ devices, which are inherently noisy in an intrusive and unavoidable way. For this reason, it is important to incorporate some understanding of this noise at every level when attempting to build solutions for practical problems using these devices. In this work, we have demonstrated the use of current quantum computers in two different paradigms to study the dynamics of real-world rail or tramway operations. More specifically, we studied the problem of (re-)scheduling a set of trains which pass through portions of a rail network that have some degree of inherent stochasticity, e.g. due to sharing infrastructure with road traffic. By focusing on this inherently noisy optimization problem, we sought to actively exploit the noise present in all current hardware as a strength.
@@
We focused on scheduling trains in a model of a portion of the Baltimore LightRail Link network, a tramway in Baltimore, Maryland, where between some stations trains must travel on tracks which share space with roadways leading to randomness in transit times. We constructed a variety of test scheduling problems starting from an actual timetable for the tramway, with the smallest problem having only a single train and the largest having twelve. These scheduling problems were easily translated to QUBOs, at which point we solved them using two types of NISQ devices. The first device employed was a quantum annealer, the D-Wave Advantage_system6.3, which has 5614 qubits and could easily fit all of our scheduling problems. The second device was the IonQ Aria-1 gate-based trapped-ion computer, a 25-qubit device which was capable of solving several of the smaller scheduling problems we constructed using QAOA. While current gate-based computers have significantly fewer qubits than quantum annealers and while QAOA is a much more expensive approach to solving QUBOs than annealing, there is significant merit in the comparison. By testing our approach on two devices with completely different computational approaches, physical architectures, and noise models we can both demonstrate the broad applicability of these techniques as well as demonstrate how variations in the implementation alter the statistics of the output.
@@
Our results show that when optimizing the schedules which govern trains passing through a stochastic zone, the effects of the random noise produce a set of schedules which when aggregated produce distributions of passing times which can potentially mimic the unpredictable delays visible in measurements of real railway traffic. Consequently, this study can serve as a starting point for the development of specialized hybrid quantum/classical algorithms that integrate NISQ solvers alongside classical approaches to aid in capturing and optimizing the stochastic behavior of railway networks. This is an especially important point as large and practically-relevant railway scheduling problems [7] are approximately 60×60$\times$60 × larger than the largest problem we have successfully solved herein. By splitting the computation between a NISQ device which handles the noisy optimization component and a classical computer which handles the larger deterministic sections of the network which encompass and connect the stochastic zone(s), it will be possible to exploit the advantages of NISQ devices while avoiding limitations due to their small size.
@@
Finally, in demonstrating the potential for NISQ devices in the realm of stochastic scheduling we have taken an important step towards a resolution of several ongoing debates within the operational research community regarding the effectiveness of QUBO formulations such as the one we present in Sec. II.2 for the planning and re-planning of railway schedules, and for more general practical applications [5] or similar problem formulations with more general QUBOs [28].
@@
Acknowledgements.
B.G acknowledges support from the National Science Center (NCN), Poland, under Project No. 2020/38/E/ST3/00269. K.D acknowledges: Scientific work co-financed from the state budget under the program of the Minister of Education and Science, Poland (pl. Polska) under the name ”Science for Society II” project number NdS-II/SP/0336/2024/01 funding amount 1000000 PLN total value of the project 1000000 PLN [Uncaptioned image]. S.D. acknowledges support the John Templeton Foundation under Grant No. 62422. We acknowledge Swiftly’s GTFS-realtime API https://swiftly.zendesk.com/hc/en-us (accessed 11-31 January 2024) for supplying real-time traffic data. K.D. acknowledges cooperation with Koleje Ślaskie sp. z o.o. (eng. Silesian Railways Ltd.) and appreciates the valuable and substantive discussions.
@@ A.1 Constraint encoding details
QUBO constraints are derived directly from the Integer Linear Programming model in Section II.1, these are.
@@
    $\sum_{t\in R_{s,j}}\left(\sum_{t^{\prime}\in R_{s^{\prime},j},t^{\prime}<t+\delta^{\text{station}}+\delta^{pass}_{s,s^{\prime}}}x_{s,j,t}x_{s^{\prime},j,t^{\prime}}\right)=0.$ (21)
@@
    $\sum_{t\in R_{s,j}}\left(\sum_{t^{\prime}\in R_{s,j^{\prime}},t-\delta^{\text{headway}}<t^{\prime}<t+\delta^{\text{headway}}}x_{s,j,t}x_{s,j^{\prime}t^{\prime}}\right)=0.$ (22)
@@
    $\sum_{t\in R_{s,j}}\left(\sum_{t^{\prime}\in R_{s,j^{\prime}},t^{\prime}<t+\delta^{\text{preparation}}+\delta^{\text{station}}}x_{s,j,t}x_{s,j^{\prime},t^{\prime}}\right)=0.$ (23)
From the definition of the time range $R_{s,j}$ from Eq. (2), we have that the outer sums in Eqs. (21)-(23) include at most $d_{\max}+1$ terms. The number of terms in the inner sums varies from $1$ to $d_{\max}+1$. If we estimate that each inner sum has an average of $\nicefrac{{(d_{\max}+1)}}{{2}}$ terms, then we expect there to be an average of $\nicefrac{{(d_{\max}+1)^{2}}}{{2}}$ terms in total for each of the three constraint types given by Eqs. (21)-(23). From the discussions in Sec. II.1 concerning the size of the sets $SP_{j}$ and $H_{s}$, we expect the number of constraints to be linear in #$\#J$ and #$\#S$. This statement relies on the reasonable assumption that there are few rolling stock constraints (i.e., that the sets $RS_{s}$ are small).
@@ A.2.1 Building the QUBO
$l^{\text{in}}_{1,{\rm PS}}$ $=19,\qquad\qquad$ $u^{\text{in}}_{1,{\rm PS}}$ $=21,$ (26)
@@
$l^{\text{in}}_{2,{\rm CS}}$ $=40,\qquad\qquad$ $u^{\text{in}}_{2,{\rm CS}}$ $=42,$ (27)
@@
From Eq. (10) there are $18$ total QUBO variables labeled $x_{s,j,t}$, where the subscripts run over the three stations $s\in\{{\rm PS},{\rm MR},{\rm CS}\}$, the two trains $j\in\{1,2\}$, and the three times possible for each variable. The ranges of the time subscripts are given by Eqs. (26) and (27), e.g., for the variables $x_{PS,1,t}$ the time index is $t\in\{19,20,21\}$.
@@
$p_{\text{sum}}\Bigg{(}\sum_{\begin{subarray}{c}t,t^{\prime}\in R_{{\rm PS},1}\\ t\neq t^{\prime}\end{subarray}}x_{\text{PS},1,t}x_{\text{PS},1,t^{\prime}}-\!\!\sum_{t\in R_{{\rm PS},1}}x^{2}_{\text{PS},1,t}\Bigg{)},$ (28)
@@
$p_{\text{pair}}\!\!\sum_{t\in R_{{\rm PS},1}}\sum_{\begin{subarray}{c}t^{\prime}\in R_{{\rm MR},1}\\ t^{\prime}<t+1+2\end{subarray}}\!\!\Big{(}x_{{\rm PS},1,t}x_{{\rm MR},1,t^{\prime}}+x_{{\rm MR},1,t^{\prime}}x_{{\rm PS},1,t}\Big{)},$ (29)
@@
$p_{\text{pair}}\!\!\sum_{t\in R_{{\rm CS},1}}\sum_{\begin{subarray}{c}t^{\prime}\in R_{{\rm CS},2}\\ t^{\prime}<t+3+1\end{subarray}}\!\!\Big{(}x_{\text{CS},1,t}x_{\text{CS},2,t^{\prime}}+x_{\text{CS},2,t^{\prime}}x_{\text{CS},1,t}\Big{)},$ (30)
@@
$\begin{split}f(\vec{x})=&\sum_{t\in R_{{\rm MR},1}}\frac{t-17}{2}x_{{\rm MR},1,t}+\sum_{t\in R_{{\rm CS},1}}\frac{t-32}{2}x_{{\rm CS},1,t}\\ +&\sum_{t\in R_{{\rm CS},2}}\frac{t-40}{2}x_{{\rm CS},2,t}+\sum_{t\in R_{{\rm MR},2}}\frac{t-55}{2}x_{{\rm MR},2,t},\end{split}$ (31)
where the vector $\vec{x}$ is shorthand for all 18 variables $x_{s,j,t}$. To translate this objective into a quadratic form suitable for a QUBO, we use the observation that $x_{s,j,t}=x^{2}_{s,j,t}$ since the variables are binary. The quadratic form $Q$ comes from combining the penalty-encoded constraints of Eqs. (28), (29), and (30) with the objective function Eq.(31). In total there are 90 non-zero elements in this formulation of our two train scheduling problem.
@@ A.2.2 Solutions
and all other variables equal to zero. The value of the objective function from Eq. (31) is $f(\vec{x})=6$. This ground state is degenerate, as delaying train $2$ by one minute at $PS$ ($x_{PS,2,59}=1\iff x_{PS,2,60}=1$) would yield the same value of the objective function and of the quadratic form.

### v2

@@ III.4 Real-world railway rescheduling perspective
For this particular problem, the input disturbed timetable includes conflicts between three pairs of trains. A train diagram representing this timetable is shown in Fig. 10a. To resolve this conflict, it is necessary that some of the trains involved in conflicts add an additional wait to their schedule. Fig. 10b shows an optimal rescheduling that resolves these conflicts, generated with a classical ILP solver. Notice that passing times of all the trains are the same (i.e., the slopes of the lines are identical). This is not what would be expected for a timetable describing the movement of trains through stochastic areas of the rail network; in such a case, the passing times should be randomly perturbed and so the slopes should vary. This feature is visible in Fig. 10c, which depict timetables corresponding to the best (lowest energy) solution returned by the D-Wave annealer and a solutions corresponding to a low-lying excited state, respectively. This observation supports the assertion that when NISQ devices are used to solve scheduling problems with this type of QUBO representation, their inherent noise may manifest in a way that can capture stochastic behavior of the underlying real-world model.
To elaborate this observation, we compare the statistics of the D-Wave output to the statistics measured from real-world rail traffic. For this comparison, we present in Fig. 11 the distribution of passing times northbound ${\rm MR}\rightarrow{\rm CS}$ and southbound ${\rm CS}\rightarrow{\rm MR}$. The data [56] was collected by aggregating passing times at peak hour during work days, as that is when the random disturbances due to road traffic are most intense.
@@ IV Conclusions
At present, quantum computing is in the era of NISQ devices, which are inherently noisy in an intrusive and unavoidable way. For this reason, it is important to incorporate some understanding of this noise at every level when attempting to build solutions for practical problems using these devices. In this work, we have demonstrated the use of current quantum computers in two different paradigms to study the dynamics of real-world rail or tramway operations. In other words, we have demonstrated the practical use of NISQ technology in solving real-world problems, particularly in optimizing transportation systems under stochastic disturbances. More specifically, we studied the problem of (re-)scheduling a set of trains which pass through portions of a rail network that have some degree of inherent stochasticity, e.g. due to sharing infrastructure with road traffic. By focusing on this inherently noisy optimization problem, we sought to actively exploit the noise present in all current hardware as a strength. Such research encourages collaboration across various fields, promotes innovation and new methodologies in both theoretical and applied research.
@@
We focused on scheduling trains in a model of a portion of the Baltimore LightRail Link network, a tramway in Baltimore, Maryland, where between some stations trains must travel on tracks which share space with roadways leading to randomness in transit times. Hereby, our research potentially improves the efficiency and reliability of public transportation systems, benefiting commuters and urban infrastructure. We constructed a variety of test scheduling problems starting from an actual timetable for the tramway, with the smallest problem having only a single train and the largest having twelve. These scheduling problems were easily translated to QUBOs, at which point we solved them using two types of NISQ devices. The first device employed was a quantum annealer, the D-Wave Advantage_system6.3, which has 5614 qubits and could easily fit all of our scheduling problems. The second device was the IonQ Aria-1 gate-based trapped-ion computer, a 25-qubit device which was capable of solving several of the smaller scheduling problems we constructed using QAOA. While current gate-based computers have significantly fewer qubits than quantum annealers and while QAOA is a much more expensive approach to solving QUBOs than annealing, there is significant merit in the comparison. By testing our approach on two devices with completely different computational approaches, physical architectures, and noise models we can both demonstrate the broad applicability of these techniques as well as demonstrate how variations in the implementation alter the statistics of the output.
@@
Our results show that when optimizing the schedules which govern trains passing through a stochastic zone, the effects of the random noise produce a set of schedules which when aggregated produce distributions of passing times which can potentially mimic the unpredictable delays visible in measurements of real railway traffic. Consequently, this study can serve as a starting point for the development of specialized hybrid quantum/classical algorithms that integrate NISQ solvers alongside classical approaches to aid in capturing and optimizing the stochastic behavior of railway networks. This is an especially important point as large and practically-relevant railway scheduling problems [7] are approximately 60×60$\times$60 × larger than the largest problem we have successfully solved herein. By splitting the computation between a NISQ device which handles the noisy optimization component and a classical computer which handles the larger deterministic sections of the network which encompass and connect the stochastic zone(s), it will be possible to exploit the advantages of NISQ devices while avoiding limitations due to their small size. The development of effective error correction strategies, such as the one assessing the D-Wave embedding strategy, or error mitigation in variational quantum algorithms such as QAOA [57] is a sound avenue for future work. Other future directions of work include the adoption of gate-based quantum optimization methods that propose better scalability in terms of the number of qubits [45, 46] or that are are not based on QUBO formulations [47].
@@
Finally, in demonstrating the potential for NISQ devices in the realm of stochastic scheduling we have taken an important step towards a resolution of several ongoing debates within the operational research community regarding the effectiveness of QUBO formulations such as the one we present in Sec. II.2 for the planning and re-planning of railway schedules, and for more general practical applications [5] or similar problem formulations with more general QUBOs [33]. To this end, our research opens avenues for further research in quantum computing applications, particularly in complex optimization scenarios and real-time system management.
@@
Data from Baltimore Light RailLink has been collected on https://github.com/iitis/Baltimore-Light-RailLink-data
Acknowledgements
B.G acknowledges support from the National Science Centre (NCN), Poland, under Project No. 2020/38/E/ST3/00269. K.D acknowledges: Scientific work co-financed from the state budget under the program of the Minister of Education and Science, Poland (pl. Polska) under the name ”Science for Society II” project number NdS-II/SP/0336/2024/01 funding amount 1000000 PLN total value of the project 1000000 PLN [Uncaptioned image]. S.D. acknowledges support the John Templeton Foundation under Grant No. 62422. We acknowledge Swiftly’s GTFS-realtime API https://swiftly.zendesk.com/hc/en-us (accessed 11-31 January 2024) for supplying real-time traffic data. K.D. acknowledges cooperation with Koleje Ślaskie sp. z o.o. (eng. Silesian Railways Ltd.) and appreciates the valuable and substantive discussions.
Author contributions statement
K.D., S.D. - conceptualization, K.D., E.D., R.R. B.G. - preparing experiments, K.D., E.D., R.R. - running experiments, K.D., E.D., R.R., B.G., S.D. - data analysis, K.D., E.D. S.D. - manuscript writing, K.D., E.D., B.G., S.D. - manuscript supervision, S.D., B.G. - funding acquisition. All authors reviewed the manuscript.
@@ A.1 Constraint encoding details
QUBO constraints are derived directly from the Integer Linear Programming model in Section II A, these are:
@@
    $\sum_{t\in R_{s,j}}\left(\sum_{t^{\prime}\in R_{s^{\prime},j},t^{\prime}<t+\delta^{\text{station}}+\delta^{pass}_{s,s^{\prime}}}x_{s,j,t}x_{s^{\prime},j,t^{\prime}}\right)=0.$ (A.1)
@@
    $\sum_{t\in R_{s,j}}\left(\sum_{t^{\prime}\in R_{s,j^{\prime}},t-\delta^{\text{headway}}<t^{\prime}<t+\delta^{\text{headway}}}x_{s,j,t}x_{s,j^{\prime}t^{\prime}}\right)=0.$ (A.2)
@@
    $\sum_{t\in R_{s,j}}\left(\sum_{t^{\prime}\in R_{s,j^{\prime}},t^{\prime}<t+\delta^{\text{preparation}}+\delta^{\text{station}}}x_{s,j,t}x_{s,j^{\prime},t^{\prime}}\right)=0.$ (A.3)
From the definition of the time range $R_{s,j}$ from Eq. (2), we have that the outer sums in Eqs. (A.1)-(A.3) include at most $d_{\max}+1$ terms. The number of terms in the inner sums varies from $1$ to $d_{\max}+1$. If we estimate that each inner sum has an average of $\nicefrac{{(d_{\max}+1)}}{{2}}$ terms, then we expect there to be an average of $\nicefrac{{(d_{\max}+1)^{2}}}{{2}}$ terms in total for each of the three constraint types given by Eqs. (A.1)-(A.3). From the discussions in Sec. II A concerning the size of the sets $SP_{j}$ and $H_{s}$, we expect the number of constraints to be linear in #$\#J$ and #$\#S$. This statement relies on the reasonable assumption that there are few rolling stock constraints (i.e., that the sets $RS_{s}$ are small).
@@ A.2.1 Building the QUBO
$l^{\text{in}}_{1,{\rm PS}}$ $=19,\qquad\qquad$ $u^{\text{in}}_{1,{\rm PS}}$ $=21,$ (A.6)
@@
$l^{\text{in}}_{2,{\rm CS}}$ $=40,\qquad\qquad$ $u^{\text{in}}_{2,{\rm CS}}$ $=42,$ (A.7)
@@
From Eq. (10) there are $18$ total QUBO variables labeled $x_{s,j,t}$, where the subscripts run over the three stations $s\in\{{\rm PS},{\rm MR},{\rm CS}\}$, the two trains $j\in\{1,2\}$, and the three times possible for each variable. The ranges of the time subscripts are given by Eqs. (A.6) and (A.7), e.g., for the variables $x_{PS,1,t}$ the time index is $t\in\{19,20,21\}$.
@@
$p_{\text{sum}}\Bigg{(}\sum_{\begin{subarray}{c}t,t^{\prime}\in R_{{\rm PS},1}\\ t\neq t^{\prime}\end{subarray}}x_{\text{PS},1,t}x_{\text{PS},1,t^{\prime}}-\!\!\sum_{t\in R_{{\rm PS},1}}x^{2}_{\text{PS},1,t}\Bigg{)},$ (A.8)
@@
$p_{\text{pair}}\!\!\sum_{t\in R_{{\rm PS},1}}\sum_{\begin{subarray}{c}t^{\prime}\in R_{{\rm MR},1}\\ t^{\prime}<t+1+2\end{subarray}}\!\!\Big{(}x_{{\rm PS},1,t}x_{{\rm MR},1,t^{\prime}}+x_{{\rm MR},1,t^{\prime}}x_{{\rm PS},1,t}\Big{)},$ (A.9)
@@
$p_{\text{pair}}\!\!\sum_{t\in R_{{\rm CS},1}}\sum_{\begin{subarray}{c}t^{\prime}\in R_{{\rm CS},2}\\ t^{\prime}<t+3+1\end{subarray}}\!\!\Big{(}x_{\text{CS},1,t}x_{\text{CS},2,t^{\prime}}+x_{\text{CS},2,t^{\prime}}x_{\text{CS},1,t}\Big{)},$ (A.10)
@@
$\begin{split}f(\vec{x})=&\sum_{t\in R_{{\rm MR},1}}\frac{t-17}{2}x_{{\rm MR},1,t}+\sum_{t\in R_{{\rm CS},1}}\frac{t-32}{2}x_{{\rm CS},1,t}\\ +&\sum_{t\in R_{{\rm CS},2}}\frac{t-40}{2}x_{{\rm CS},2,t}+\sum_{t\in R_{{\rm MR},2}}\frac{t-55}{2}x_{{\rm MR},2,t},\end{split}$ (A.11)
where the vector $\vec{x}$ is shorthand for all 18 variables $x_{s,j,t}$. To translate this objective into a quadratic form suitable for a QUBO, we use the observation that $x_{s,j,t}=x^{2}_{s,j,t}$ since the variables are binary. The quadratic form $Q$ comes from combining the penalty-encoded constraints of Eqs. (A.8), (A.9), and (A.10) with the objective function Eq.(A.11). In total there are 90 non-zero elements in this formulation of our two train scheduling problem.
@@ A.2.2 Solutions
and all other variables equal to zero. The value of the objective function from Eq. (A.11) is $f(\vec{x})=6$. This ground state is degenerate, as delaying train $2$ by one minute at $PS$ ($x_{PS,2,59}=1\iff x_{PS,2,60}=1$) would yield the same value of the objective function and of the quadratic form.
@@
A.2.3 Spectral Properties
A complete and concrete specification of the QUBO to be minimized, formed from the combination of the constraints from Eqs. (A.8), (A.9), and (A.10) alongside the objective from Eq. (A.11), requires that specific values be chosen for $p_{\rm sum}$ and $p_{\rm pair}$. As discussed in Section 2.2, identifying good choices for these parameters can be challenging. In this work, we have chosen to consider each problem instance with two choices of penalty values, one “overlapping” case with lower penalties chosen such that the sets of feasible and infeasible solutions overlap in energy and one “split” case with higher penalties such that there is a moderate energy gap separating the two sets of solutions. Figure A.1 shows the distributions of energies for these two choices applied to the 18-variable QUBO discussed in this section. On a broad, coarse-grained level the two distributions look essentially the same up to an overall rescaling of the energies, but the histogram of just a small window of energies around the optimal solution reveals the difference. With the smaller parameters, there is overlap between the sets of feasible and infeasible solutions and there are no large gaps (small gaps appear as all energies are integer multiples of $1/2$). With larger penalty parameters, the distribution of energies reveals clusters of solutions with similar energies separated by large gaps. These clusters correspond to different constraint violations, hence all feasible solutions are isolated in their own cluster at the lowest values of the objective.
For the larger problems we consider it rapidly becomes intractable to exhaustively enumerate the collection of solution energies, but as is clear from the results obtained with the D-Wave annealer on an 11 train, 182 variable problem shown Fig. 3 this dichotomy between overlapping and split spectra are present in all the problem instances we consider. Note that the same comb-like structure as appears in Fig. 3b and Fig. 3d would appear in the distribution of all energies in Fig. A.1 if the histograms used finer binning, as in the zoomed histograms. When considering all energies a coarse binning is necessary for the feasible solution set to be visible – the D-Wave annealer is by design biased toward sampling low-energy and feasible solutions.
Figure A.1: Histograms of QUBO values for the possible solutions to the 2 train, 18 variable problem with two choices of constraint parameters. The top panels include all $2^{18}$ possible solutions, and make it clear that there is little difference in the large-scale structure of the two choices beyond a rescaling of the energy. The bottom panels show only the feasible solutions as well as the infeasible solutions closest to optimal. With the lower choice of penalties, the two sets of solutions overlap, whereas with the larger penalties, the two sets of solutions are entirely separate.